Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 588     Entry time: 19 Dec 2018, 13:17
Author: Adriano Violante 
Location: General 
Severity: Information 
Subject: Status report after dispatch meeting 09.30 

SPB/SFX

  • SFX part1 cabling preparation ongoing (WCBs mounting and labeling). Good progress.
  • CRLs will be brought out of XTD9 tomorrow.
  • Scaffolding below D.23 has been removed. An official communication that now the area is completely accessible is still lacking.

FXE

SQS

  • For the loops 9-13 the HW is ready. CAS projects have to be prepared.
  • Some issues with middle layer devices have been notified to CAS (Cyril)

SCS

MID

  • Karabo 2.3.2 has been deployed apart from the AGIPD control (waiting for tests by DET group before update).
  • CY is now back working on the Mpots (SASE2 and 3)

HED

OPT Lasers

XRO

​​XPD

​​VAC

AE

  • SASE1 power was back yesterday. SPB and FXE have been brought up again yesterday afternoon. Today morning because of unplanned and badly communicated MPS works the power was cut again ==> additional, unplanned efforts needed.
  • Hazem will be OCD during Christmas, only for PLCs. 
  • No uTCA OCD during Christmas.

EET

ITDM

DET

CAS

  • 2.3.2 Deployed in SASE2
  • SPB update postponed to January
  • SASE1 planned for today
  • No OCD until 09.01.

AoB

  • Next Disptach meeting on 07.01.2019 at 9:30
ELOG V3.1.4-7c3fd00